OpenStudy (anonymous):

hey y'all me again :p So ... at a high school, the average grade for the english 10 provincial is 64, with a standard deviation of 10. If 20 students scored between 73 and 86 on the exam, how many students took the exam? Use z-score in your solution. thanks :)

jimthompson5910 (jim_thompson5910):

First convert each raw score to a z-score So convert x = 73 to a z-score z = (x - mu)/(sigma) z = (73 - 64)/(10) z = 0.9 and then convert x = 86 to a z-score z = (x - mu)/(sigma) z = (86 - 64)/(10) z = 2.2 So the raw scores of 73 and 86 convert to the z-scores 0.9 and 2.2 respectively

jimthompson5910 (jim_thompson5910):

You're then told that "20 students scored between 73 and 86 on the exam" So 20 students have a z-score between 0.9 and 2.2

jimthompson5910 (jim_thompson5910):

Your task is to find the area under the standard normal curve from z = 0.9 to z = 2.2 Let's call this area p. If x students total took the test, then xp students scored between z = 0.9 and z = 2.2 Put another way, if x students total took the test, then xp students scored between 73 and 86

jimthompson5910 (jim_thompson5910):

But we know that 20 students scored between 73 and 86, so xp = 20

jimthompson5910 (jim_thompson5910):

So all you need to do is find p. Once you have p, solve for x to get your answer.

OpenStudy (anonymous):

is p then .170 ?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

okay so 118 students took the test ?

jimthompson5910 (jim_thompson5910):

yes if you round up

OpenStudy (anonymous):

thank u so so much

OpenStudy (anonymous):

i was so lost, thanks

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

how did you find that p equaled 170?